OpenStudy (anonymous):

Explain two ways you could evaluate (f + g)(2). Explain two ways you could evaluate (f + g)(2). @Mathematics

OpenStudy (psymon):

well, you could do f(2), get an answer, then add it to g(2). Or you can combine all the like terms and combine f(x) and g(x), then plug in 2 and solve. Both get the same answer. Ex. \[f(x)=3x ^{2}+4\] \[g(x) = 5x ^{2}-6x+1\] If I evaluate f(2) I get 3(2)^2 + 4 = 16 If I evaluate g(2) I get 5(2)^2 - 6(2) + 1 = 9. Add the two together andI get 25. Now if I do it the other way and combine the two functions first. \[f(x)+g(x) = 8x ^{2}-6x+5\] Now I'll evaluate (f+g)(2) 8(2)^2 - 6(2) + 5 = 32 - 12 + 5 = 25. So you can see both give the same answer :3

OpenStudy (anonymous):

i need help with another problem

OpenStudy (anonymous):

Does (f - g)(x) result in the same function as (g - f)(x)?

OpenStudy (uri):

(f + g)(2) f(2)+g(2) I find this the easiest ^

OpenStudy (uri):

Yes @aliballa1 It's same.

OpenStudy (psymon):

Wouldn't they just be opposites? One would be just the negative of the other.

OpenStudy (uri):

Oh yea..

OpenStudy (anonymous):

What is a good strategy for being successful with word problems?